Find the equation of the line shown. Enter your answer in point-slope form.
Slope:3/-4, Ordered pair (-10.5,-8)

Answer by Boreal(15235)

point slope formula is y-y1=m(x-x1), m slope (x1, y1) point
this is y+8=(-3/4)(x+10.5)